OpenStudy (anonymous):

Find the general form of the equation of the plane witht he given characteristics: Passes through (2,2,1) and (-1,1,-1) and is perpendicular to 2x-3y+z=3

OpenStudy (anonymous):

plane vector: (2-(-1),2-1,1-(-1)) = (3,1,2) =a perpendicular to 2x-3y+z=3 means perpendicular to (2,-3,1) = b and a c = axb = (7,1,-12) (x stands for cross product) so Plane equation is: 7x+y-12z = 0

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

@LivyLou

OpenStudy (anonymous):

Where did you get the (3,1,2) and (7,1,-12) from?

OpenStudy (anonymous):

(3,1,2) is a vector that goes from (-1,1,-1) to (2,2,1) (7,1,-12) is a vector product of a and b

OpenStudy (anonymous):

Does using the vector product give the perpendicular vector?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay thanks